# A General Control-Theoretic Approach for Reinforcement Learning: Theory and Algorithms

### v1

@@ C.2 Lunar Lander (Continuous) under CBRL LQR
Figure 8 presents our numerical results for the three state-of-the-art baselines (continuous actions) and our CBRL approach, each run over five independent random seeds. All parameters in (11) are initialized uniformly within $(0,1)$ and then learned using our control-policy-parameter gradient ascent iteration (5). We consider four sets of initial parameters (see Table 2) to validate the robustness of our CBRL approach. Figure 8(a) and 8(b) clearly demonstrate that our CBRL approach outperforms all other baselines w.r.t. both the number of episodes and running time. Figure 8(c) – 8(f) shows the learning behavior of CBRL parameters given different initialization.
@@
(a) Return v.s. Episode
(b) Return v.s. Time
@@
Figure 8: Learning curves of LunarLanderContinuous-v2 over five independent runs. The solid line shows the mean and the shaded area depicts the standard deviation. (a) and (b): Return vs. number of episodes and running time, respectively, for our CBRL approach in comparison with the Linear policy, PPO, and DDPG. (c) – (f): Learning behavior of CBRL parameters, initialized by Table 2.
C.3 Mountain Car (Continuous) under CBRL Piecewise-LQR
@@
As depicted in Figure 9, Mountain Car (Continuous) consists of a car placed in a valley with the goal of accelerating the car to reach the goal on top of the right hill by applying a force on the car to the left or right. The state of the system is $x=[s,v]$ in terms of the position of the car $s$ and the velocity of the car $v$. With random initialization at the bottom of the valley, the problem is considered “solved” upon achieving an average return of $90$.
@@
Recall that the LQR controller is not sufficient to solve the Mountain Car (Continuous) problem, even if all the parameters of the system are known (e.g., mass of the car and gravity), because a nonlinear controller is required. Consequently, we consider a piecewise-LQR controller that partitions the state space into two regions: LQR 1 and LQR 2 (see Figure 9). The target state $x^{*}=[s^{*},v^{*}]$ for LQR 1 and LQR 2 are respectively selected as $x^{*}=[-1.2,0]$ and $x^{*}=[0.6,0]$, where $-1.2$ and $0.6$ represent the position of the left hill and the right hill. We address the problem within the context of our CBRL approach by exploiting the following general matrix form for the piecewise-LQR dynamics
@@
which solely takes into account general physical relationships and laws with the unknown parameters $a_{0},a_{1},b_{0},c_{0}$ to be learned, where $e=x-x^{*}$ and we select $v^{*}=0,\dot{v}^{*}=0$, and $s^{*}=-1.2$ or $0.6$.
@@
Figure 10 presents our numerical results for the three state-of-the-art baselines (continuous actions) and our CBRL approach, each run over five independent random seeds. All parameters in (12) are initialized uniformly within $(0,1)$ and then learned using our control-policy-parameter gradient ascent iteration (5). We consider four sets of initial parameters (see Table 3) to validate the robustness of our CBRL approach. Figure 10(a) and 10(b) clearly demonstrate that our CBRL approach outperforms all other baselines w.r.t. both the number of episodes and running time, in addition to demonstrating a more stable training process. Figure 10(c) – 10(f) shows the learning behavior of CBRL parameters given different initialization.
@@
(a) Return v.s. Episode
(b) Return v.s. Time
@@
Figure 10: Learning curves of MountainCarContinuous-v0 over five independent runs. The solid line shows the mean and the shaded area depicts the standard deviation. (a) and (b): Return vs. number of episodes and running time, respectively, for our CBRL approach in comparison with the Linear policy, PPO, and DDPG. (c) – (f): Learning behavior of CBRL parameters, initialized by Table 3.
C.4 Pendulum under CBRL Piecewise-LQR
@@
As depicted in Figure 11, Pendulum consists of a link attached at one end to a fixed point and the other end being free, with the goal of swinging up to an upright position by applying a torque on the free end. The state of the system is $x=[\theta,\dot{\theta}]$ in terms of the angle of the link $\theta$ and the angular velocity of the link $\dot{\theta}$. With random initial position of the link, each episode comprises $200$ steps and the problem is considered “solved” upon achieving an average return of $-200$.
@@
which solely takes into account general physical relationships and laws with the unknown parameters $a_{0},a_{1},b_{0},c_{0},d_{0}$ to be learned, where $e=x-x^{*}$ and we select $\ddot{\theta}^{*}=0$.
@@
Figure 12 presents our numerical results for the three state-of-the-art baselines (continuous actions) and our CBRL approach, each run over five independent random seeds. All parameters in (13) are initialized uniformly within $(0,1)$ and then learned using our control-policy-parameter gradient ascent iteration (5). We consider four sets of initial parameters (see Table 4) to validate the robustness of our CBRL approach. Figure 12(a) illustrates a comparable performance between our CBRL approach and the best baseline DDPG in terms of the convergence w.r.t. the number of episodes, while Figure 12(b) clearly demonstrates that CBRL outperforms all other baselines w.r.t. running time, in addition to demonstrating a more stable training process. Figure 12(c) – 12(f) shows the learning behavior of CBRL parameters given different initialization.
@@
(a) Return v.s. Episode
(b) Return v.s. Time
@@
Figure 12: Learning curves of Pendulum-v1 over five independent runs. The solid line shows the mean and the shaded area depicts the standard deviation. (a) and (b): Return vs. number of episodes and running time, respectively, for our CBRL approach in comparison with the Linear policy, PPO, and DDPG. (c) – (f): Learning behavior of CBRL parameters, initialized by Table 4.
C.5 Implementation
@@
As previously noted, the evaluation of our CBRL approach includes comparisons against the three baselines of DQN for discrete actions, DDPG for continuous actions and PPO, which are selected as the state-of-the-art RL algorithms for solving the control tasks under consideration. The additional baseline of replacing the nonlinear policy of PPO with a linear policy was again included because the optimal policy for some problems, e.g., Cart Pole, is known to be linear. This Linear algorithm solely replaces the policy network of PPO with a linear policy while maintaining all other components unchanged. More specifically, we decrease the number of hidden layers in the policy network by $1$, reduce the hidden layer size of the policy network to $16$, and remove all (nonlinear) activation functions in the policy network.
Stepsize.
The stepsize selection is an important implementation issue, just like in nonconvex optimization and RL. We address this issue by using similar methods employed in standard RL, i.e., the adaptive stepsize selection in our implementation of Algorithm 1. More concretely, we simply reduce the stepsize $\eta$ in (5) by a factor of $0.99$ each time achieving the “solved” return.
Hyperparameters

### v2

@@ Appendix B Experimental Results
Table 3 and Figure 8 present numerical results for the three state-of-the-art baselines (continuous actions) and our CBRL approach, each run over five independent random seeds. All parameters in (16) are initialized uniformly within $(0,1)$ and then learned using our control-policy-parameter gradient ascent iteration (6). We consider four sets of initial parameters (see Table 4) to validate the robustness of our CBRL approach. Figure 8(c) – 8(f) illustrates the learning behavior of CBRL parameters given different initialization.
Table 3 and Figure 8(a) and 8(b) clearly demonstrate that our CBRL approach provides far superior performance, in terms of both mean and standard deviation, over all baselines w.r.t. both the number of episodes and running time, in addition to demonstrating a more stable training process. More specifically, our CBRL approach outperforms all baselines with a significant improvement in mean return over independent random environment runs from Gymnasium (Towers et al. 2023), together with a significant reduction in the standard deviation of the return over these independent runs, thus rendering a more robust solution approach where the performance of each run is much closer to the mean than under the corresponding baseline results. In fact, the relative percentage difference [^2] For the relative comparison of improved mean return performance of our CBRL method over the next-best RL method, since some of the performance values are negative and $|M_{RL}|<M_{CBRL}$, we use the standard relative percentage difference formula $\frac{M_{CBRL}-M_{RL}}{|M_{RL}|}\times 100$. Similarly, for the relative comparison of reduced standard deviation of return performance from the next-best RL method to our CBRL method, since all the performance values are positive, we use the standard relative percentage difference formula $\frac{S_{RL}-S_{CBRL}}{S_{CBRL}}\times 100$. Refer to https://en.wikipedia.org/wiki/Relative˙change in improved mean return performance of our CBRL approach over the next-best RL method DDPG (based on mean performance at $500$ episodes) is $309\%$ and $110\%$ at $200$ and $500$ episodes, respectively; and the relative percentage difference in reduced standard deviation of return performance from the next-best RL method DDPG to our CBRL approach is $115\%$ and $288\%$ at $200$ and $500$ episodes, respectively. Moreover, the relative percentage difference in improved mean return performance of our CBRL approach over the PPO RL method is $361\%$ and $390\%$ at $200$ and $500$ episodes, respectively; and the relative percentage difference in reduced standard deviation of return performance from the PPO RL method to our CBRL approach is $639\%$ and $559\%$ at $200$ and $500$ episodes, respectively. We note that the baseline algorithms often crash and terminate sooner than the more successful landings of our CBRL approach, resulting in the significantly worse performance exhibited in Figure 2, Table 3 and Figure 8.
We also note an important difference between Figure 2(a),(b) and Figure 8(a),(b), namely the shaded areas in Figure 2(a),(b) represent one form of variability across independent random seeds with the same parameter initialization, whereas the shaded areas in Figure 8(a),(b) represent the combination of two forms of variability – one across independent random seeds with the same parameter initialization and the other across different random parameter initializations.
Table 3: Mean and Standard Deviation of Return of RL Methods for LunarLanderContinuous-v2
Episode CBRL Linear PPO DDPG
Number Mean Std. Dev. Mean Std. Dev. Mean Std. Dev. Mean Std. Dev.
50 $-128.81$ $276.58$ $-531.05$ $115.63$ $-403.56$ $143.06$ $-337.22$ $98.64$
100 $101.33$ $262.09$ $-533.6$ $130.04$ $-260.08$ $238.84$ $-145.05$ $42.29$
150 $258.84$ $31.52$ $-510.13$ $123.56$ $-188.50$ $121.78$ $-154.48$ $43.2$
200 $279.24$ $18.12$ $-513.44$ $129.48$ $-106.87$ $133.97$ $-133.51$ $38.96$
250 $280.36$ $17.19$ $-475.43$ $125.81$ $-127.29$ $81.82$ $-143.46$ $112.20$
300 $279.19$ $18.42$ $-473.77$ $127.78$ $-109.72$ $79.65$ $-73.1$ $70.24$
350 $280.08$ $17.14$ $-479.01$ $124.67$ $-45.12$ $140.02$ $-9.04$ $98.49$
400 $279.51$ $17.67$ $-458.10$ $131.61$ $-75.52$ $111.54$ $70.04$ $120.93$
450 $278.96$ $19.22$ $-445.52$ $143.65$ $-91.77$ $105.02$ $51.15$ $99.21$
500 $280.09$ $17.16$ $-492.67$ $126.01$ $-96.50$ $113.10$ $133.20$ $66.54$
@@
(a) Return vs. Episode
(b) Return vs. Time
@@
Figure 8: Learning curves of LunarLanderContinuous-v2 over five independent runs. The solid line shows the mean and the shaded area depicts the standard deviation. (a) and (b): Return vs. number of episodes and running time, respectively, for our CBRL approach (over the five independent runs and the four initializations in Table 4) in comparison with the Linear policy, PPO, and DDPG (over the five independent runs). (c) – (f): Learning behavior of CBRL parameters, initialized by Table 4.
Mountain Car (Continuous) under CBRL Piecewise-LQR
@@
As depicted in Figure 9 and described in the main body of the paper, Mountain Car (Continuous) consists of a car placed in a valley with the goal of accelerating the car to reach the target at the top of the hill on the right by applying a force on the car to the left or right. The state of the system is given by $x=[s,v]$ in terms of the position of the car $s$ and the velocity of the car $v$.
@@
Recall that the LQR controller is not sufficient to solve the Mountain Car (Continuous) problem, even if all the parameters of the system are known (e.g., mass of the car and gravity), because a nonlinear controller is required. Consequently, we consider a piecewise-LQR controller that partitions the state space into two regions: LQR 1 and LQR 2 (see Figure 9). The target state $x^{*}=[s^{*},v^{*}]$ for LQR 1 and LQR 2 are respectively selected as $x^{*}=[-1.2,0]$ and $x^{*}=[0.6,0]$, where $-1.2$ and $0.6$ correspondingly represent the position of the left hill and the right hill. We address the problem within the context of our CBRL approach by exploiting the following general matrix form for the piecewise-LQR dynamics
@@
which solely takes into account general physical relationships and laws, with the unknown parameters $a_{0},a_{1},b_{0},c_{0}$ to be learned where $e=x-x^{*}$ and we select $v^{*}=0,\dot{v}^{*}=0$, and $s^{*}=-1.2$ or $0.6$.
@@
Table 5 and Figure 10 present numerical results for the three state-of-the-art baselines (continuous actions) and our CBRL approach, each run over five independent random seeds. All parameters in (17) are initialized uniformly within $(0,1)$ and then learned using our control-policy-parameter gradient ascent iteration (6). We consider four sets of initial parameters (see Table 6) to validate the robustness of our CBRL approach. Figure 10(c) – 10(f) illustrates the learning behavior of CBRL parameters given different initialization.
Table 5 and Figure 10(a) and 10(b) clearly demonstrate that our CBRL approach provides superior performance, in terms of both mean and standard deviation, over all baselines w.r.t. both the number of episodes and running time, in addition to demonstrating a more stable training process. More specifically, our CBRL approach outperforms all baselines with a significant improvement in mean return over independent random environment runs from Gymnasium (Towers et al. 2023), together with a significant reduction in the standard deviation of the return over these independent runs, thus rendering a more robust solution approach where the performance of each run is much closer to the mean than under the corresponding baseline results. In fact, the relative percentage difference [^3] For the relative comparison of improved mean return performance of our CBRL method over the next-best RL method, since all the performance values are positive, we use the standard relative percentage difference formula $\frac{M_{CBRL}-M_{RL}}{M_{RL}}\times 100$. Similarly, for the relative comparison of reduced standard deviation of return performance from the next-best RL method to our CBRL method, since all the performance values are positive, we use the standard relative percentage difference formula $\frac{S_{RL}-S_{CBRL}}{S_{CBRL}}\times 100$. Refer to https://en.wikipedia.org/wiki/Relative˙change in improved mean return performance of our CBRL approach over the next-best RL method DDPG (based on mean performance at $500$ episodes) is $0.9\%$ and $1\%$ at $200$ and $500$ episodes, respectively; and the relative percentage difference in reduced standard deviation of return performance from the next-best RL method DDPG to our CBRL approach is $324\%$ and $257\%$ at $200$ and $500$ episodes, respectively. Moreover, the relative percentage difference in improved mean return performance of our CBRL approach over the PPO RL method is $41\%$ and $23\%$ at $200$ and $500$ episodes, respectively; and the relative percentage difference in reduced standard deviation of return performance from the PPO RL method to our CBRL approach is $5743\%$ and $729\%$ at $200$ and $500$ episodes, respectively.
We note an important difference between Figure 3(a),(b) and Figure 10(a),(b), namely the shaded areas in Figure 3(a),(b) represent one form of variability across independent random seeds with the same parameter initialization, whereas the shaded areas in Figure 10(a),(b) represent the combination of two forms of variability – one across independent random seeds with the same parameter initialization and the other across different random parameter initializations.
Table 5: Mean and Standard Deviation of Return of RL Methods for MountainCarContinuous-v0
Episode CBRL Linear PPO DDPG
Number Mean Std. Dev. Mean Std. Dev. Mean Std. Dev. Mean Std. Dev.
50 $79.34$ $1.02$ $61.09$ $16.84$ $59.62$ $7.32$ $74.20$ $7.20$
100 $93.03$ $0.37$ $61.64$ $17.19$ $65.01$ $9.69$ $89.10$ $4.78$
150 $93.63$ $0.21$ $63.12$ $19.53$ $69.58$ $10.57$ $91.98$ $1.47$
200 $93.63$ $0.21$ $61.94$ $21.83$ $66.36$ $12.27$ $92.82$ $0.89$
250 $93.63$ $0.21$ $62.30$ $25.59$ $71.35$ $13.15$ $93.18$ $0.41$
300 $93.63$ $0.21$ $62.00$ $25.12$ $71.25$ $12.27$ $93.62$ $0.35$
350 $93.63$ $0.21$ $57.13$ $34.55$ $76.23$ $3.03$ $93.39$ $0.51$
400 $93.63$ $0.21$ $62.42$ $25.91$ $81.32$ $1.63$ $93.15$ $0.81$
450 $93.63$ $0.21$ $66.19$ $20.54$ $81.10$ $2.02$ $92.67$ $0.85$
500 $93.63$ $0.21$ $61.52$ $28.84$ $76.31$ $1.74$ $92.72$ $0.75$
@@
(a) Return vs. Episode
(b) Return vs. Time
@@
Figure 10: Learning curves of MountainCarContinuous-v0 over five independent runs. The solid line shows the mean and the shaded area depicts the standard deviation. (a) and (b): Return vs. number of episodes and running time, respectively, for our CBRL approach (over the five independent runs and the four initializations in Table 6) in comparison with the Linear policy, PPO, and DDPG (over the five independent runs). (c) – (f): Learning behavior of CBRL parameters, initialized by Table 6.
Pendulum under CBRL Piecewise-LQR
@@
As depicted in Figure 11 and described in the main body of the paper, Pendulum consists of a link attached at one end to a fixed point and the other end being free, with the goal of swinging up to an upright position by applying a torque on the free end. The state of the system is given by $x=[\theta,\dot{\theta}]$ in terms of the angle of the link $\theta$ and the angular velocity of the link $\dot{\theta}$.
@@
which solely takes into account general physical relationships and laws, with the unknown parameters $a_{0},a_{1},b_{0},c_{0},d_{0}$ to be learned where $e=x-x^{*}$ and we select $\ddot{\theta}^{*}=0$.
@@
Table 7 and Figure 12 present numerical results for the three state-of-the-art baselines (continuous actions) and our CBRL approach, each run over five independent random seeds. All parameters in (18) are initialized uniformly within $(0,1)$ and then learned using our control-policy-parameter gradient ascent iteration (6). We consider four sets of initial parameters (see Table 8) to validate the robustness of our CBRL approach. Figure 12(c) – 12(f) illustrates the learning behavior of CBRL parameters given different initialization.
Table 7 and Figure 12(a) and 12(b) clearly demonstrate that our CBRL approach provides far superior performance, in terms of both mean and standard deviation, over all baselines w.r.t. both the number of episodes and running time upon convergence of our CBRL algorithm after a relatively small number of episodes, in addition to demonstrating a more stable training process. More specifically, even with only four partitions for the difficult nonlinear Pendulum RL task, our CBRL approach outperforms all baselines after $150$ episodes with a significant improvement in mean return over independent random environment runs from Gymnasium (Towers et al. 2023), together with a significant reduction in the standard deviation of the return over these independent runs, thus rendering a more robust solution approach where the performance of each run is much closer to the mean than under the corresponding baseline results. In fact, the relative percentage difference [^4] For the relative comparison of improved mean return performance of our CBRL method over the next-best RL method, since all the performance values are negative, we use the reversed standard relative percentage difference formula $\frac{|M_{RL}-M_{CBRL}|}{|M_{CBRL}|}\times 100$. Similarly, for the relative comparison of reduced standard deviation of return performance from the next-best RL method to our CBRL method, since all the performance values are positive, we use the standard relative percentage difference formula $\frac{S_{RL}-S_{CBRL}}{S_{CBRL}}\times 100$. Refer to https://en.wikipedia.org/wiki/Relative˙change in improved mean return performance of our CBRL approach over the next-best RL method DDPG (based on mean performance at $500$ episodes) is $40\%$ and $53\%$ at $200$ and $500$ episodes, respectively; and the relative percentage difference in reduced standard deviation of return performance from the next-best RL method DDPG to our CBRL approach is $210\%$ and $263\%$ at $200$ and $500$ episodes, respectively. Moreover, the relative percentage difference in improved mean return performance of our CBRL approach over the PPO RL method is $197\%$ and $168\%$ at $200$ and $500$ episodes, respectively; and the relative percentage difference in reduced standard deviation of return performance from the PPO RL method to our CBRL approach is $275\%$ and $793\%$ at $200$ and $500$ episodes, respectively.
We note an important difference between Figure 4(a),(b) and Figure 12(a),(b), namely the shaded areas in Figure 4(a),(b) represent one form of variability across independent random seeds with the same parameter initialization, whereas the shaded areas in Figure 12(a),(b) represent the combination of two forms of variability – one across independent random seeds with the same parameter initialization and the other across different random parameter initializations.
Table 7: Mean and Standard Deviation of Return of RL Methods for Pendulum-v1
Episode CBRL Linear PPO DDPG
Number Mean Std. Dev. Mean Std. Dev. Mean Std. Dev. Mean Std. Dev.
50 $-1378.78$ $43.89$ $-1331.73$ $36.25$ $-1147.35$ $181.20$ $-942.18$ $170.83$
100 $-1004.75$ $101.66$ $-1145.30$ $53.57$ $-757.11$ $205.66$ $-355.09$ $324.16$
150 $-284.26$ $130.17$ $-1088.67$ $83.73$ $-559.65$ $151.39$ $-241.16$ $162.85$
200 $-172.48$ $48.32$ $-1062.06$ $102.63$ $-511.85$ $181.28$ $-240.94$ $149.57$
250 $-152.80$ $37.43$ $-1058.88$ $92.55$ $-561.78$ $384.15$ $-222.18$ $151.71$
300 $-145.56$ $44.16$ $-1057.02$ $90.29$ $-524.10$ $259.46$ $-221.78$ $150.40$
350 $-141.80$ $36.63$ $-1036.65$ $88.37$ $-325.66$ $133.77$ $-222.83$ $146.68$
400 $-149.34$ $51.70$ $-1044.34$ $100.90$ $-218.37$ $100.83$ $-220.28$ $148.80$
450 $-139.65$ $32.36$ $-1032.82$ $87.11$ $-296.94$ $184.32$ $-220.50$ $148.68$
500 $-143.95$ $40.91$ $-999.41$ $78.75$ $-385.50$ $365.39$ $-220.38$ $148.54$
@@
(a) Return vs. Episode
(b) Return vs. Time
@@
Figure 12: Learning curves of Pendulum-v1 over five independent runs. The solid line shows the mean and the shaded area depicts the standard deviation. (a) and (b): Return vs. number of episodes and running time, respectively, for our CBRL approach (over the five independent runs and the four initializations in Table 8) in comparison with the Linear policy, PPO, and DDPG (over the five independent runs). (c) – (f): Learning behavior of CBRL parameters, initialized by Table 8.
Implementation
@@
As previously noted, the evaluation of our CBRL approach includes comparisons against the three baselines of DQN for discrete actions, DDPG for continuous actions and PPO, which are selected as the state-of-the-art RL algorithms for solving the RL tasks under consideration. The additional baseline of replacing the nonlinear policy of PPO with a linear policy was again included because the optimal policy for some problems, e.g., Cart Pole, is known to be linear. This Linear algorithm solely replaces the policy network of PPO with a linear policy while maintaining all other components unchanged. More specifically, we decrease the number of hidden layers in the policy network by $1$, reduce the hidden layer size of the policy network to $16$, and remove all (nonlinear) activation functions in the policy network.
Hyperparameters.
@@
Computing Infrastructure.
All experiments in this work were implemented and executed on an 8G-memory Macbook Air with a M1 Chip.